What is the total entropy change for the decomposition of hydrogen peroxide.


2H2O2(l) ---> 2H2O(l) + O2(g)


Given


H2O2(l)

(J/mol x K) = 109.6

(kJ/mol) = -187.8


H2O(l)

(J/mol x K) = 69.9

(kJ/mol) = -285.8


O2(g)

(J/mol x K) = 205.0

(kJ/mol) = 0


1
Expert's answer
2022-04-26T04:02:04-0400

At 298K as a standard temperature:

S°(O2, g) = 205.0 J mol−1 K−1

S°(H2O, l) = 69.9 J mol−1 K−1

S°(H2O2, l) = 109.6 J mol−1 K−1


Solution:

The entropy change in a chemical reaction is given by the sum of the entropies of the products minus the sum of the entropies of the reactants.

ΔS° =∑nS°(products) − ∑mS°(reactants)


Balanced chemical equation:

2H2O2(l) → 2H2O(l) + O2(g)


From the balanced equation we can write the equation for ΔS° (the change in the standard molar entropy for the reaction):

ΔS° = S°(O2, g) + 2×S°(H2O, l) − 2×S°(H2O2, l)

ΔS° = 205.0 + 2×69.9 − 2×109.6 = 125.6 (J mol−1 K−1)

ΔS° = 125.6 J mol−1 K−1


Answer: The total entropy change for the decomposition of hydrogen peroxide is 125.6 J mol−1 K−1